OpenStudy (anonymous):

Suppose it takes 2 and a 1/2 hours to drive from your hours to the lake at 48mi/h . How long will your return trip take at 40mi/h ?

OpenStudy (a_clan):

48*(5/2)=40*X

OpenStudy (anonymous):

2.5*48=120 miles 120/40=3 hours